OpenStudy (swag):

The distance between the points (3,6) and (0,10) is equal to 5. True False

OpenStudy (anonymous):

False

OpenStudy (phi):

use the distance formula http://www.purplemath.com/modules/distform.htm for your problem \[ d= \sqrt{ (3-0)^2 + (6-10)^2 }\] that simplifies to \[ d= \sqrt{ 3^2 + (-4)^2 }\] have you ever heard of a 3-4-5 right triangle ?

OpenStudy (swag):

Yes I think I have.

OpenStudy (phi):

d= sqrt(9+16)= sqrt(25)= 5